OpenStudy (anonymous):

what about Given the points: A(0,0), B (6,3), C(5,5), and D(-1,2). What two segments must be parallel? Choose all possible answers. AB || DC AB || BC AB || AD DC || AD BC || AD BC || CD

OpenStudy (anonymous):

anyone know the answer to that

OpenStudy (tkhunny):

Have you considered calculating the slope of each segment? \(\overline{AB}\) \(\overline{AC}\) \(\overline{AD}\) \(\overline{BC}\) \(\overline{BD}\) \(\overline{CD}\)

OpenStudy (anonymous):

No cause I don't know how to Can you please tell me the answer tkhunny

OpenStudy (anonymous):

I don't know this at all

OpenStudy (anonymous):

Plz help me

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to a segment created by joining points (1, 0) and (5, -2) and also passes through the midpoint of that segement. Any form is acceptable.

OpenStudy (tkhunny):

A(0,0), B (6,3), C(5,5), D(-1,2) \(Slope \overline{AB} = \dfrac{3-0}{6-0}\) \(Slope \overline{AC} = \dfrac{5-0}{5-0}\) \(Slope \overline{AD} = \dfrac{-1-0}{2-0}\) You do the other three. Don't look for answers. Look for understanding.